Support department cost allocation—reciprocal services method

Davis Snowflake & Co. produces Christmas stockings in its Cutting and Sewing departments. The Maintenance and Security departments support the production of the stockings. Costs from the Maintenance Department are allocated based on machine hours, and costs from the Security Department are allocated based on asset value. Information about each department is provided in the following table:

Maintenance
Department
Security
Department
Cutting
Department
Sewing
Department
Machine hours 800    2,000 6,800    11,200
Asset value $2,000    $1,420 $3,000 $5,000
Department cost $38,160    $16,960 $67,000 $85,000

Determine the total cost of each production department after allocating all support department costs to the production departments using the reciprocal services method.

Cutting
Department
Sewing
Department
Production departmentsʼ total costs $ $

Support Department Cost Allocation—Reciprocal Services Method

Blue Africa Inc. produces laptops and desktop computers. The company’s production activities mainly occur in what the company calls its Laser and Forming departments. The Cafeteria and Security departments support the company’s production activities and allocate costs based on the number of employees and square feet, respectively. The total cost of the Security Department is $243,000. The total cost of the Cafeteria Department is $674,000. The number of employees and the square footage in each department are as follows:

Employees Square Feet
Security Department 10        550       
Cafeteria Department 20        2,400       
Laser Department 40        4,800       
Forming Department 50        800       

Using the reciprocal services method of support department cost allocation, determine the total costs from the Security Department that should be allocated to the Cafeteria Department and to each of the production departments.

Cafeteria
Department
Laser
Department
Forming
Department
Security Department cost allocation $ $ $
Answer #1
Cutting Sewing
Production Department Total Cost $           87,776 $     1,19,344
Workings:
To
From Maintenance Security Cutting Sewing
Service Departmentl costs $           38,160 $        16,960 $       67,000 $         85,000
Maintenance (3:4:3) $         -42,400 $           4,240 $       14,416 $         23,744
Cafeteria (6:2:2) $             4,240 $       -21,200 $         6,360 $         10,600
Total Cost $                    -   $                  -   $       87,776 $     1,19,344
Let Total Maintenance department cost be x and Security Department be y
x = 38160 + (2000/10000 * y)
y = 16960 + (2000/20000 * x)
By solving above:
y = 16960 + (2000/20000 * x)
y = 16960 + (2000/20000 * (38160 + (2000/10000 * y)))
y = $21200
also,
x = 38160 + (2000/10000 * y)
x = 38160 + (2000/10000 * 21200)
x = $42400
Cafeteria Department Laser Department Forming Department
Security Departmentl cost allocation $        96,000 $     1,92,000 $          32,000
Workings:
Let Total Security department cost be x and of Cafeteria be y
x = 243000 + (10/100 * y)
y = $674000 + (2400/8000 * x)
By solving above:
y = $674000 + (2400/8000 * x)
y = $674000 + (2400/8000 *( 243000 + (10/100 * y)))
y = $770000
also,
x = 243000 + (10/100 * y)
x = 243000 + (10/100 * 770000)
x = $320000
Security Departmentl cost allocation
Cafeteria Department = [$320000 * (2400/8000)] = $        96,000
Laser Department = [$320000 * (4800/8000)] = $     1,92,000
Forming Department = [$320000 * (800/8000)] = $        32,000